The economy is at the natural level of unemployment of 6% and is described by the following equations: 1. oyt (a) How does a 2 year long disinflation from 8% to 4% inflation look like? (b) How does a 4 year disinflation from 8% to 4% look like when the Central Bank commits to lower inflation?Please show all your work
Answer #1

Answers

a) From Phillips Curve(PC):

\pi_t - \pi_t_-_1 = -2(\mu_t - \mu_n )

and from given Okun's Law(OL) we have: \mu_t - \mu_t_-_1 = -0.4(g_y_t - 0.03)

For Year 1,

Given \pi_t = 4 %, \pi_t^e= 8 % and  \mu_n = 6 %

Based on PC, \mu_t =\frac{\pi_t - \pi_t_-1-2\mu_n}{2} = 8%

Now, substituting \mu_t value in OL equation we will get g_y_t

-0.04g_y_t = \mu_t - \mu_t_-_1 - 0.012

-0.04g_y_t = 0.08 - 0.06 - 0.012

-0.04g_y_t = 0.008

g_y_t = -0.2

Similarly, we will solve for 2nd-year

b) The value for inflation, expected inflation, the natural rate of unemployment is given.

We will calculate first \mu_t based on PC equation. Thereafter we will calculate in g_y_t the similar fashion as done above.

Output Shortfall as mentioned in notes: (g_y_t - g_n_t) = 2.5(\mu_t - \mu_n)

Substituting the values we will derive g_n_t here. Therafter the output shortfall can be calculated.